# Acylindrical hyperbolicity and the centers of Artin groups that are not free of infinity

### v2

@@ 3. Reduced clique-cube complexes and actions on them
(1) The assertion is clear from parts (1), (2), and (3) of Lemma 4.2.
(2) Parts (1), (2), (3) and (4) of Lemma 4.2 and Lemma 4.3 imply that the sequence specified in the assertion is a sequence of separating hyperplanes from $J_{i_{1},0}$ to $J_{i_{r+1},2(\mathfrak{m}+r)n+1}$. Note that $i_{1}=1$ and $i_{r+1}=1$ by definition.
We show that the sequence contains a hyperplane of $v_{i}$-type for any $i\in\{1,\ldots,k\}$ and $v_{i}\in V(\Gamma_{i})$. Recall that $(V_{i_{1}},\dots,V_{i_{r}},V_{i_{r+1}})$ (4.1) is a closed path on the tree $T$. Thus, if an edge $(V_{i},V_{j})\in E(T)$ is contained in the closed path, then so is the inverse edge $(V_{j},V_{i})$. In addition, because $T$ is a spanning tree of $Q(\Gamma)$ and the closed path $(V_{i_{1}},\dots,V_{i_{r}},V_{i_{r+1}})$ passes through every vertex at least once, any $i\in\{1,\ldots,k\}$, $V_{i}$ is contained in the closed path as a vertex. Additionally, recall that $T$ has the root $V_{1}$ and that $i<j$ only if $V_{i}$ is not farther than $V_{j}$ from $V_{1}$ in $T$.
@@
First, suppose that $i=1$. Note that $i_{1}=1$ and set $j=i_{2}$. Take $a\in\{1,2,\ldots,r\}$ such that $a\neq 1$, $i_{a}=1$, and $i_{a-1}=j$. Then, the sequence in the assertion contains the two subsequences
$J_{i_{1},2\mathfrak{m}n+1},J_{i_{1},2\mathfrak{m}n+2},\ldots,J_{i_{1},2\mathfrak{m}n+2l(i_{1},i_{2})-1},$
which are of $v_{1,1}$-type, $v_{1,1}$-type, $\ldots$, $v_{1,l(1,j)}$-type, and
$J_{i_{a},2\mathfrak{m}n+2((a-2)n+l(i_{a-1},i_{a}))},J_{i_{a},2\mathfrak{m}n+2((a-2)n+l(i_{a-1},i_{a}))+1},\ldots,J_{i_{a},2\mathfrak{m}n+2(a-1)n},$
which are of $v_{1,l(j,1)}$-type, $v_{1,l(j,1)+1}$-type, $\ldots$, $v_{1,n}$-type. Note that $l(j,1)=l(1,j)$ by Lemma 4.1. Take any vertex $v_{1}\in V(\Gamma_{1})$. Then, there exists some $l\in\{1,\ldots,n\}$ such that $v_{1,l}=v_{1}$ by Lemma 4.1. Hence, the sequence in the assertion contains a hyperplane of $v_{1}$-type.
Next, suppose that $i\neq 1$. Take the smallest $a\in\{1,\ldots,r\}$ such that $i_{a}=i$. Because $i_{1}=1$ and $i\neq 1$, we have $a\neq 1$. We set $j=i_{a-1}$. Then, we have $a^{\prime}\in\{1,\ldots,r\}$ such that $a\leq a^{\prime}$, $i_{a^{\prime}}=i$, and $i_{a^{\prime}+1}=j$. The sequence contains the two subsequences
$J_{i_{a},2\mathfrak{m}n+2((a-2)n+l(i_{a-1},i_{a}))},J_{i_{a},2\mathfrak{m}n+2((a-2)n+l(i_{a-1},i_{a}))+1},\ldots,J_{i_{a},2\mathfrak{m}n+2(a-1)n},$
which are of $v_{i,l(j,i)}$-type, $v_{i,l(j,i)+1}$-type, $\ldots$, $v_{i,n}$-type, and
$J_{i_{a^{\prime}},2\mathfrak{m}n+2(a^{\prime}-1)n+1},J_{i_{a^{\prime}},2\mathfrak{m}n+2((a^{\prime}-1)n+1)},\ldots,J_{i_{a^{\prime}},2\mathfrak{m}n+2((a^{\prime}-1)n+l(i_{a^{\prime}},i_{a^{\prime}+1}))-1},$
which are of $v_{i,1}$-type, $v_{i,1}$-type, $\ldots$, $v_{i,l(j,i)}$-type. Note that $l(j,i)=l(i,j)$ by Lemma 4.1. Take any vertex $v_{i}\in V(\Gamma_{i})$. Then, there exists some $l\in\{1,\ldots,n\}$ such that $v_{i,l}=v_{i}$ by Lemma 4.1. Hence, the sequence in the assertion contains a hyperplane of $v_{i}$-type. ∎
@@
We show $(1)\Rightarrow(2)$ in Theorem 1.2. Consider $\gamma\in A_{\Gamma}$ defined by (4.9) and hyperplanes
$J:=J_{1,0}\text{ and }J^{\prime}:=J_{1,2(\mathfrak{m}+r)n+1},$
which are of $v_{1,n}$-type and $v_{1,1}$-type, respectively. We will confirm conditions (i), (ii), and (iii) in Theorem 2.9.
(i) $\gamma$ skewers $(J,J^{\prime})$. Indeed, it is clear that
$J^{+}\supsetneq(\gamma^{-1}(J^{\prime+})\supsetneq\gamma(J^{+})\supsetneq)J^{\prime+}\supsetneq\gamma^{2}(J^{+}).$
(ii) We show that $J$ and $J^{\prime}$ are strongly separated. Take any hyperplane $H$ with $J\cap H\neq\emptyset$. When $H$ is of $v_{i}$-type for some $i$ and $v_{i}\in V(\Gamma_{i})$, take a hyperplane $H^{\prime}$ of $v_{i}$-type separating $A_{V_{0}}$ and $\gamma A_{V_{0}}$ such that $J\cap H^{\prime}=\emptyset$ by part (2) of Lemma 4.5. Then, $H\cap H^{\prime}=\emptyset$ by Remark 2.13. Because $J\subset H^{\prime-},H^{\prime+}\supset J^{\prime}$, $J\cap H\neq\emptyset$ and $H\cap H^{\prime}=\emptyset$, we have $J^{\prime}\cap H=\emptyset$.
(iii) We show $Stab(J)\cap Stab(J^{\prime})=\{1\}$. We devide the proof into three parts: (iii-1) we show that $Stab(J)\cap Stab(J^{\prime})\subset A_{V_{0}}$; (iii-2) we show that $Stab(J)\cap Stab(J^{\prime})\subset\gamma A_{V_{0}}\gamma^{-1}$; (iii-3) we show that $Stab(J)\cap Stab(J^{\prime})=\{1\}$ using (iii-1) and (iii-2).
(iii-1) Note that for any $i\in\{1,\ldots,k\}$ and any $v_{i}\in V(\Gamma_{i})$, we have at least one sequence of separating hyperplanes $P^{\prime}_{1},\ldots,P^{\prime}_{M^{\prime}}$ from $A_{V_{0}}$ to $\gamma A_{V_{0}}$ such that $P^{\prime}_{1}$ is of $v_{i}$-type and $P^{\prime}_{M^{\prime}}$ is of $v_{1,n}$-type. For example, we can take such a sequence by considering a subsequence of the sequence in part (2) of Lemma 4.5. For any $i\in\{1,\ldots,k\}$ and any $v_{i}\in V(\Gamma_{i})$, we can take a longest sequence of separating hyperplanes $P_{1},\ldots,P_{M}$ from $A_{V_{0}}$ to $\gamma A_{V_{0}}$ such that $P_{1}$ is of $v_{i}$-type and $P_{M}$ is of $v_{1,n}$-type, where
$A_{V_{0}}\in P_{1}^{-},P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\ni\gamma A_{V_{0}}$
by taking decompositions by appropriate connected components
@@
Note that $P_{1},\ldots,P_{M}\in\{J_{j,d}\}_{j\in\{1,\ldots,k\},d\in\{1,\ldots,2(\mathfrak{m}+r)n\}}$ by Lemma 4.4. By noting $P_{M}\in\{J_{1,d}\}_{d\in\{1,\ldots,2(\mathfrak{m}+r)n\}}$ and part (1) of Lemma 4.5 for the case $i=1$, we have $P_{M}^{+}\supsetneq J^{\prime+}$.
Now assume that $Stab(J)\cap Stab(J^{\prime})\not\subset A_{V_{0}}$. Take $g\in Stab(J)\cap Stab(J^{\prime})$ with $g\not\in A_{V_{0}}$. Note that $g^{-1}\in Stab(J)\cap Stab(J^{\prime})$. Then, we have a hyperplane $H$ of $v_{i}$-type for some $i\in\{1,\ldots,k\}$ and some $v_{i}\in V(\Gamma_{i})$, separating $A_{V_{0}}$ and $gA_{V_{0}}$. We take a longest sequence of separating hyperplanes $P_{1},\ldots,P_{M}$ from $A_{V_{0}}$ to $\gamma A_{V_{0}}$ such that $P_{1}$ is of $v_{i}$-type and $P_{M}$ is of $v_{1,n}$-type. Then we have
$g\gamma A_{V_{0}},g^{-1}\gamma A_{V_{0}}\in P_{M}^{+}$
by $P_{M}^{+}\supsetneq J^{\prime+}$. We take a connected component $H^{-}$ of $C_{\Gamma,\Gamma_{0}}\!\setminus\!\!\setminus H$ such that $A_{V_{0}}t\in H^{-}$. Then the other connected component $H^{+}$ satisfies $gA_{V_{0}}\in H^{+}$. Because $H\cap J\neq\emptyset$, we have $H\cap J^{\prime}=\emptyset$ by (ii). Thus $H$ can not separate $\gamma A_{V_{0}}$ and $g\gamma A_{V_{0}}$. Hence we have either
  1. (iii-1-a)
    $A_{V_{0}},\gamma A_{V_{0}},g\gamma A_{V_{0}}\in H^{-}\text{ and }gA_{V_{0}}\in H^{+}$ or
  2. (iii-1-b)
    $A_{V_{0}}\in H^{-}\text{ and }gA_{V_{0}},\gamma A_{V_{0}},g\gamma A_{V_{0}}\in H^{+}$.
Assume that the case (iii-1-a) occurs. Then $H$ does not separate $A_{V_{0}}$ and $\gamma A_{V_{0}}$ and thus $H\notin\{J_{i,d}\}_{d\in\{1,\ldots,2(\mathfrak{m}+r)n\}}$ by Lemma 4.4. Hence $H$ and $P_{1}$ are different hyperplanes of the same $v_{i}$-type, and thus they can not intersect by Remark 2.13. Then we have either $H^{-}\supsetneq P_{1}^{+}$ or $H^{+}\supsetneq P_{1}^{+}$. However $H^{+}\supsetneq P_{1}^{+}$ can not occur because $\gamma A_{V_{0}}\notin H^{+}$ and $\gamma A_{V_{0}}\in P_{1}^{+}$. Therefore
@@
$gA_{V_{0}}\in H^{+},H^{-}\supsetneq P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\ni g\gamma A_{V_{0}}.$
Then, $Q_{1}=H,Q_{2}=P_{1},\ldots,Q_{M+1}=P_{M}$ is a sequence of separating hyperplanes from $gA_{V_{0}}$ to $g\gamma A_{V_{0}}$ such that $Q_{1}$ is of $v_{i}$-type and $Q_{M+1}$ is of $v_{1,n}$-type. Thus, $g^{-1}Q_{1},g^{-1}Q_{2},\ldots,g^{-1}Q_{M+1}$ is a sequence of separating hyperplanes from $A_{V_{0}}$ to $\gamma A_{V_{0}}$ such that $g^{-1}Q_{1}$ is of $v_{i}$-type and $g^{-1}Q_{M+1}$ is of $v_{1,n}$-type, which contradicts the fact that the sequence $P_{1},\ldots,P_{M}$ is longest.
Next, assume that the case (iii-1-b) occurs. Then $H$ does not separate $gA_{V_{0}}$ and $g\gamma A_{V_{0}}$, that is, $g^{-1}H$ does not separate $A_{V_{0}}$ and $\gamma A_{V_{0}}$ and thus $g^{-1}H\notin\{J_{i,d}\}_{d\in\{1,\ldots,2(\mathfrak{m}+r)n\}}$ by Lemma 4.4. Hence $g^{-1}H$ and $P_{1}$ are different hyperplanes of the same $v_{i}$-type, and thus they can not intersect by Remark 2.13. Then we have either $g^{-1}H^{+}\supsetneq P_{1}^{+}$ or $g^{-1}H^{-}\supsetneq P_{1}^{+}$. However $g^{-1}H^{-}\supsetneq P_{1}^{+}$ can not occur because $\gamma A_{V_{0}}\notin g^{-1}H^{-}$ and $\gamma A_{V_{0}}\in P_{1}^{+}$. Therefore
@@
$g^{-1}A_{V_{0}}\in g^{-1}H^{-},g^{-1}H^{+}\supsetneq P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\ni g^{-1}\gamma A_{V_{0}}.$
Then, $Q_{1}=H,Q_{2}=gP_{1},\ldots,Q_{M+1}=gP_{M}$ is a sequence of separating hyperplanes from $A_{V_{0}}$ to $\gamma A_{V_{0}}$ such that $Q_{1}$ is of $v_{i}$-type and $Q_{M+1}$ is of $v_{1,n}$-type, which contradicts the fact that the sequence $P_{1},\ldots,P_{M}$ is longest.
@@
(iii-2) We show $Stab(J)\cap Stab(J^{\prime})\subset\gamma A_{V_{0}}\gamma^{-1}$. This can be shown by a similar way with (iii-1). Indeed we can lead a contradiction by assuming $Stab(J)\cap Stab(J^{\prime})\not\subset\gamma A_{V_{0}}\gamma^{-1}$.
@@
Suppose that $g\in A_{V_{0}\setminus\{w^{\prime}\in V_{0}\mid w^{\prime}<w\}}=A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\geq w\}}$. Let $\ell$ be a path from $w_{0}=A_{V_{0}}$ to $w_{2(\mathfrak{m}+r)n}=\gamma A_{V_{0}}$ that diagonally penetrates each of the cubes $K_{1},\ldots,K_{2(\mathfrak{m}+r)n}$ in order. We note that the set of all hyperplanes intersecting the path $\ell$ is $\{J_{i,d}\}_{i\in\{1,\ldots,k\},d\in\{1,\ldots,2(\mathfrak{m}+r)n\}}$. Therefore, $\ell$ contains a vertex
(4.10) $\gamma_{\flat}(1)\cdots\gamma_{\flat}(\mathfrak{j}(w)-1)\lambda_{1}\cdots\lambda_{l(w_{\mathfrak{j}(w)})-1}\lambda_{l(w_{\mathfrak{j}(w)})}(\mathfrak{j}(w))A_{V_{0}},$

### v3

@@ 3. Reduced clique-cube complexes and actions on them
(1) The assertion is clear from parts (1), (2), (3) and (4) of Lemma 4.4.
(2) Parts (1), (2), (3) and (4) of Lemma 4.4 and Lemma 4.5 imply that sequences in (2-i), (2-ii) and (2-iii-$a$) are sequences of separating hyperplanes.
(3) The sequence in the assertion is a sequence of separating hyperplanes. We need to show that the sequence contains a hyperplane of $v_{i}$-type for any $i\in\{1,\ldots,k\}$ and any $v_{i}\in V_{i}$. Recall that $(V_{i_{1}},\dots,V_{i_{r}},V_{i_{r+1}})$ (4.6) is a closed path on the tree $T$. Thus, if an edge $(V_{i},V_{j})\in E(T)$ is contained in the closed path, then so is the inverse edge $(V_{j},V_{i})$. In addition, because $T$ is a spanning tree of $Q(\Gamma)$ and the closed path $(V_{i_{1}},\dots,V_{i_{r}},V_{i_{r+1}})$ passes through every vertex at least once, for any $i\in\{1,\ldots,k\}$, $V_{i}$ is contained in the closed path as a vertex. Additionally, recall that $T$ has the root $V_{1}$ and that $i<j$ only if $V_{i}$ is not farther than $V_{j}$ from $V_{1}$ in $T$.
@@
First, suppose that $i=1$. Note that $i_{1}=1$ and set $j=i_{2}$. Take $a\in\{1,2,\ldots,r\}$ such that $a\neq 1$, $i_{a}=1$, and $i_{a-1}=j$. Then, the sequence in the assertion contains the first half of the sequence of (2-iii-$1$):
$J_{i_{1},2(1+\mathfrak{m})n+1},J_{i_{1},2(1+\mathfrak{m})n+2},\ldots,J_{i_{1},2(1+\mathfrak{m})n+2l_{\flat}(i_{1},i_{2})-1},$
which are of $v_{1,1}$-type, $v_{1,1}$-type, $\ldots$, $v_{1,l_{\flat}(1,j)}$-type, and the second half of the sequence of (2-iii-$(a-1)$):
$J_{i_{a},2(1+\mathfrak{m})n+2((a-2)n+l_{\flat}(i_{a-1},i_{a}))},J_{i_{a},2(1+\mathfrak{m})n+2((a-2)n+l_{\flat}(i_{a-1},i_{a}))+1},\ldots,$
$J_{i_{a},2(1+\mathfrak{m})n+2(a-1)n},$
which are of $v_{1,l_{\flat}(j,1)}$-type, $v_{1,l_{\flat}(j,1)+1}$-type, $\ldots$, $v_{1,n}$-type. Note that $l_{\flat}(j,1)=l_{\flat}(1,j)$ by Lemma 4.2. Take any vertex $v_{1}\in V_{1}$. Then, there exists some $l\in\{1,\ldots,n\}$ such that $v_{1,l}=v_{1}$ by Lemma 4.2. Hence, the sequence in the assertion contains a hyperplane of $v_{1}$-type.
Next, suppose that $i\neq 1$. Take the smallest $a\in\{1,\ldots,r\}$ such that $i_{a}=i$. Because $i_{1}=1$ and $i\neq 1$, we have $a\neq 1$. We set $j=i_{a-1}$. Then, we have $a^{\prime}\in\{1,\ldots,r\}$ such that $a\leq a^{\prime}$, $i_{a^{\prime}}=i$, and $i_{a^{\prime}+1}=j$. The sequence in the assertion contains the second half of the sequence of (2-iii-$(a-1)$):
$J_{i_{a},2(1+\mathfrak{m})n+2((a-2)n+l_{\flat}(i_{a-1},i_{a}))},J_{i_{a},2(1+\mathfrak{m})n+2((a-2)n+l_{\flat}(i_{a-1},i_{a}))+1},\ldots,$
$J_{i_{a},2(1+\mathfrak{m})n+2(a-1)n},$
which are of $v_{i,l_{\flat}(j,i)}$-type, $v_{i,l_{\flat}(j,i)+1}$-type, $\ldots$, $v_{i,n}$-type, and the first half of the sequence of (2-iii-$a^{\prime}$):
$J_{i_{a^{\prime}},2(1+\mathfrak{m})n+2(a^{\prime}-1)n+1},J_{i_{a^{\prime}},2(1+\mathfrak{m})n+2((a^{\prime}-1)n+1)},\ldots,$
$J_{i_{a^{\prime}},2(1+\mathfrak{m})n+2((a^{\prime}-1)n+l_{\flat}(i_{a^{\prime}},i_{a^{\prime}+1}))-1},$
which are of $v_{i,1}$-type, $v_{i,1}$-type, $\ldots$, $v_{i,l_{\flat}(j,i)}$-type. Note that $l_{\flat}(j,i)=l_{\flat}(i,j)$ by Lemma 4.2. Take any vertex $v_{i}\in V_{i}$. Then, there exists some $l\in\{1,\ldots,n\}$ such that $v_{i,l}=v_{i}$ by Lemma 4.2. Hence, the sequence in the assertion contains a hyperplane of $v_{i}$-type. ∎
@@
We show $(1)\Rightarrow(2)$ in Theorem 4.1. Consider $\gamma_{\natural\sharp\flat}\in A_{\Gamma}$ defined by (4.28) and hyperplanes
$J:=J_{1,0}\text{ and }J^{\prime}:=J_{1,2(1+\mathfrak{m}+r)n+1},$
which are of $v_{1,n}$-type and $v_{1,1}$-type, respectively. We now verify conditions (i), (ii) and (iii) in Theorem 2.9.
(i) $\gamma_{\natural\sharp\flat}$ skewers $(J,J^{\prime})$. Indeed, it is clear that
$J^{+}\supsetneq\gamma_{\natural\sharp\flat}^{-1}(J^{\prime+})\supsetneq\gamma_{\natural\sharp\flat}(J^{+})\supsetneq J^{\prime+}\supsetneq\gamma_{\natural\sharp\flat}^{2}(J^{+}).$
(ii) We show that $J$ and $J^{\prime}$ are strongly separated. Take any hyperplane $H$ with $J\cap H\neq\emptyset$. When $H$ is of $v_{i}$-type for some $i$ and $v_{i}\in V_{i}$, take a hyperplane $H^{\prime}$ of $v_{i}$-type separating $A_{V_{0}}$ and $\gamma_{\natural\sharp\flat}A_{V_{0}}$ such that $J\cap H^{\prime}=\emptyset$ by part (3) of Lemma 4.7. Then, $H\cap H^{\prime}=\emptyset$ by Remark 2.14. Because $J\subset H^{\prime-},H^{\prime+}\supset J^{\prime}$, $J\cap H\neq\emptyset$ and $H\cap H^{\prime}=\emptyset$, we have $J^{\prime}\cap H=\emptyset$.
(iii) We show $Stab(J)\cap Stab(J^{\prime})=\{1\}$. We divide the proof into three parts: (iii-1) we show that $Stab(J)\cap Stab(J^{\prime})\subset A_{V_{0}}$; (iii-2) we show that $Stab(J)\cap Stab(J^{\prime})\subset\gamma_{\natural\sharp\flat}A_{V_{0}}\gamma_{\natural\sharp\flat}^{-1}$; (iii-3) we show that $Stab(J)\cap Stab(J^{\prime})=\{1\}$ using (iii-1) and (iii-2).
(iii-1) Note that for any $i\in\{1,\ldots,k\}$ and any $v_{i}\in V_{i}$, we have at least one sequence of separating hyperplanes $P^{\prime}_{1},\ldots,P^{\prime}_{M^{\prime}}$ from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$ such that $P^{\prime}_{1}$ is of $v_{i}$-type and $P^{\prime}_{M^{\prime}}$ is of $v_{1,n}$-type. For example, we can take such a sequence by considering a subsequence of the sequence in part (3) of Lemma 4.7. For any $i\in\{1,\ldots,k\}$ and any $v_{i}\in V_{i}$, we can take a longest sequence of separating hyperplanes $P_{1},\ldots,P_{M}$ from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$ such that $P_{1}$ is of $v_{i}$-type and $P_{M}$ is of $v_{1,n}$-type, where
$A_{V_{0}}\in P_{1}^{-},P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\ni\gamma_{\natural\sharp\flat}A_{V_{0}}$
by taking decompositions into appropriate connected components
@@
Note that $P_{1},\ldots,P_{M}\in\{J_{j,d}\}_{j\in\{1,\ldots,k\},d\in\{1,\ldots,2(1+\mathfrak{m}+r)n\}}$ by Lemma 4.6. By noting $P_{M}\in\{J_{1,d}\}_{d\in\{1,\ldots,2(1+\mathfrak{m}+r)n\}}$ and applying part (1) of Lemma 4.7 for the case $i=1$, we have $P_{M}^{+}\supsetneq J^{\prime+}$.
Now assume that $Stab(J)\cap Stab(J^{\prime})\not\subset A_{V_{0}}$. Take $g\in Stab(J)\cap Stab(J^{\prime})$ with $g\not\in A_{V_{0}}$. Note that $g^{-1}\in Stab(J)\cap Stab(J^{\prime})$. Then, we have a hyperplane $H$ of $v_{i}$-type for some $i\in\{1,\ldots,k\}$ and some $v_{i}\in V_{i}$, separating $A_{V_{0}}$ and $gA_{V_{0}}$. We take a longest sequence of separating hyperplanes $P_{1},\ldots,P_{M}$ from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$ such that $P_{1}$ is of $v_{i}$-type and $P_{M}$ is of $v_{1,n}$-type. Then we have
$g\gamma_{\natural\sharp\flat}A_{V_{0}},g^{-1}\gamma_{\natural\sharp\flat}A_{V_{0}}\in P_{M}^{+}$
by $P_{M}^{+}\supsetneq J^{\prime+}$. We take a connected component $H^{-}$ of $C_{\Gamma,\Gamma_{0}}\!\setminus\!\!\setminus H$ such that $A_{V_{0}}\in H^{-}$. Then the other connected component $H^{+}$ satisfies $gA_{V_{0}}\in H^{+}$. Because $H\cap J\neq\emptyset$, we have $H\cap J^{\prime}=\emptyset$ by (ii). Thus, $H$ cannot separate $\gamma_{\natural\sharp\flat}A_{V_{0}}$ and $g\gamma_{\natural\sharp\flat}A_{V_{0}}$. Hence we have either
(iii-1-$a$) $A_{V_{0}},\gamma_{\natural\sharp\flat}A_{V_{0}},g\gamma_{\natural\sharp\flat}A_{V_{0}}\in H^{-}\text{ and }gA_{V_{0}}\in H^{+}$ or
(iii-1-$b$) $A_{V_{0}}\in H^{-}\text{ and }gA_{V_{0}},\gamma_{\natural\sharp\flat}A_{V_{0}},g\gamma_{\natural\sharp\flat}A_{V_{0}}\in H^{+}$.
Assume that case (iii-1-$a$) occurs. Then $H$ does not separate $A_{V_{0}}$ and $\gamma_{\natural\sharp\flat}A_{V_{0}}$ and thus $H\notin\{J_{i,d}\}_{d\in\{1,\ldots,2(1+\mathfrak{m}+r)n\}}$ by Lemma 4.6. Hence, $H$ and $P_{1}$ are different hyperplanes of the same $v_{i}$-type, and thus they cannot intersect by Remark 2.14. Then we have either $H^{-}\supsetneq P_{1}^{+}$ or $H^{+}\supsetneq P_{1}^{+}$. However, $H^{+}\supsetneq P_{1}^{+}$ cannot occur because $\gamma_{\natural\sharp\flat}A_{V_{0}}\notin H^{+}$ and $\gamma_{\natural\sharp\flat}A_{V_{0}}\in P_{1}^{+}$. Therefore,
@@
$gA_{V_{0}}\in H^{+},H^{-}\supsetneq P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\ni g\gamma_{\natural\sharp\flat}A_{V_{0}}.$
Then, $Q_{1}=H,Q_{2}=P_{1},\ldots,Q_{M+1}=P_{M}$ is a sequence of separating hyperplanes from $gA_{V_{0}}$ to $g\gamma_{\natural\sharp\flat}A_{V_{0}}$ such that $Q_{1}$ is of $v_{i}$-type and $Q_{M+1}$ is of $v_{1,n}$-type. Thus, $g^{-1}Q_{1},g^{-1}Q_{2},\ldots,g^{-1}Q_{M+1}$ is a sequence of separating hyperplanes from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$ such that $g^{-1}Q_{1}$ is of $v_{i}$-type and $g^{-1}Q_{M+1}$ is of $v_{1,n}$-type, which contradicts the fact that the sequence $P_{1},\ldots,P_{M}$ is the longest.
Next, assume that the case (iii-1-$b$) occurs. Then $H$ does not separate $gA_{V_{0}}$ and $g\gamma_{\natural\sharp\flat}A_{V_{0}}$; that is, $g^{-1}H$ does not separate $A_{V_{0}}$ and $\gamma_{\natural\sharp\flat}A_{V_{0}}$ and thus $g^{-1}H\notin\{J_{i,d}\}$ $({d\in\{1,\ldots,2(1+\mathfrak{m}+r)n\}})$ by Lemma 4.6. Hence, $g^{-1}H$ and $P_{1}$ are different hyperplanes of the same $v_{i}$-type, and thus they cannot intersect by Remark 2.14. Then we have either $g^{-1}H^{+}\supsetneq P_{1}^{+}$ or $g^{-1}H^{-}\supsetneq P_{1}^{+}$. However $g^{-1}H^{-}\supsetneq P_{1}^{+}$ can not occur because $\gamma_{\natural\sharp\flat}A_{V_{0}}\notin g^{-1}H^{-}$ and $\gamma_{\natural\sharp\flat}A_{V_{0}}\in P_{1}^{+}$. Therefore
@@
$g^{-1}A_{V_{0}}\in g^{-1}H^{-},g^{-1}H^{+}\supsetneq P_{1}^{+}\supsetneq P_{2}^{+}\supsetneq\cdots\supsetneq P_{M-1}^{+}\supsetneq P_{M}^{+}\ni g^{-1}\gamma_{\natural\sharp\flat}A_{V_{0}}.$
Then, $Q_{1}=H,Q_{2}=gP_{1},\ldots,Q_{M+1}=gP_{M}$ is a sequence of separating hyperplanes from $A_{V_{0}}$ to $\gamma_{\natural\sharp\flat}A_{V_{0}}$ such that $Q_{1}$ is of $v_{i}$-type and $Q_{M+1}$ is of $v_{1,n}$-type, which contradicts the fact that the sequence $P_{1},\ldots,P_{M}$ is the longest.
@@
(iii-2) We show $Stab(J)\cap Stab(J^{\prime})\subset\gamma_{\natural\sharp\flat}A_{V_{0}}\gamma_{\natural\sharp\flat}^{-1}$. This can be shown in a similar way as (iii-1). Indeed we reach a contradiction by assuming $Stab(J)\cap Stab(J^{\prime})\not\subset\gamma_{\natural\sharp\flat}A_{V_{0}}\gamma_{\natural\sharp\flat}^{-1}$.
@@
Suppose that $g\in A_{V_{0}\setminus\{w^{\prime}\in V_{0}\mid w^{\prime}\prec w\}}=A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succeq w\}}$. We show that $g\in A_{\{w^{\prime\prime}\in V_{0}\mid w^{\prime\prime}\succ w\}}$.
Figure 10. Cubes penetrated by $\ell$ when $k=3$ and $l_{\sharp}(\mathfrak{j})=1$. In this case, $g$ fixes cubes $[A_{V_{0}},A_{U_{1}}]$ and $[v_{1,1}v_{2,1}v_{3,1}A_{V_{0}},A_{U_{1}}]$, which are penetrated by $\ell$. Then $g$ fixes three edges $[A_{V_{0}\sqcup\{v_{1,1},v_{2,1}\}},A_{U_{1}}]$, $[A_{V_{0}\sqcup\{v_{1,1},v_{3,1}\}},A_{U_{1}}]$ and $[v_{1,1}A_{V_{0}\sqcup\{v_{2,1},v_{3,1}\}},A_{U_{1}}]$, and thus three squares $[A_{V_{0}\sqcup\{v_{1,1}\}},A_{U_{1}}]$, $[v_{1,1}A_{V_{0}\sqcup\{v_{2,1}\}},A_{U_{1}}]$ and $[v_{1,1}A_{V_{0}\sqcup\{v_{3,1}\}},A_{U_{1}}]$. Hence $g$ fixes the cube $[v_{1,1}A_{V_{0}},A_{U_{1}}]$. By a similar argument, $g$ fixes the cube $[v_{1,1}v_{2,1}A_{V_{0}},A_{U_{1}}]$. Therefore, $g$ fixes not only $A_{V_{0}}$ and $v_{1,1}v_{2,1}v_{3,1}A_{V_{0}}$, but also $v_{1,1}A_{V_{0}}$ and $v_{1,1}v_{2,1}A_{V_{0}}$.
(iii-3-$a$) We consider the case where $h=1$; that is, $w\in W(1)$. Then we have $v_{i_{\sharp}(w),l_{\sharp}(w)}\in V_{\ast}$. We consider a vertex on $\ell$ of the form
(4.35) $\lambda_{1}\cdots\lambda_{l_{\sharp}(w)-1}\lambda_{l_{\sharp}(w)}A_{V_{0}},$